# KPC-cF: Aspect-Based Sentiment Analysis via Implicit-Feature Alignment with Corpus Filtering

### v4

@@ Abstract
Investigations into Aspect-Based Sentiment Analysis (ABSA) for Korean industrial reviews are notably lacking in the existing literature. Our research proposes an intuitive and effective framework for ABSA in low-resource languages such as Korean. It optimizes prediction labels by integrating translated benchmark and unlabeled Korean data. Using a model fine-tuned on translated data, we pseudo-labeled the actual Korean NLI set. Subsequently, we applied LaBSE and MSP-based filtering to this pseudo-NLI set as implicit feature, enhancing Aspect Category Detection and Polarity determination through additional training. Incorporating dual filtering, this model bridged dataset gaps, achieving positive results in Korean ABSA with minimal resources. Through additional data injection pipelines, our approach aims to utilize high-resource data and construct effective models within communities, whether corporate or individual, in low-resource language countries. Compared to English ABSA, our framework showed an approximately 3% difference in F1 scores and accuracy. We release the dataset and our code for Korean ABSA, at this link111https://github.com/namkibeom/KPC-cF.
KPC-cF: Aspect-Based Sentiment Analysis via
Implicit-Feature
Alignment with Corpus Filtering
Kibeom Nam Department of Data Science, Hongik University KorABSA Laboratory, MODULABS skarlqja68@gmail.com
@@ 1 Introduction
In low-resource downstream tasks such as Korean ABSA, constraints exist in constructing ABSA systems that are socially and industrially beneficial (e.g., obtaining accurate labels and high-quality training data, building a efficient serving model). Addressing this challenge is fundamentally crucial for the practical implementation of multilingual ABSA leveraging the advantages of language models Zhang et al. (2021); Lin et al. (2023). On the other hand, ABSA utilizing Large Language Models like ChatGPT can perform labeling through prompt tuning. However, it still has limitations compared to small-scale models in terms of classifier metrics and resource for training and inference Wang et al. (2023); Wu et al. (2023).
@@
In this study, we derive pseudo-labels for real Korean reviews using machine-translated English ABSA data, inspired by the past research Balahur and Turchi (2012); Hoshino et al. (2024). Moreover, we employ Dual filtering on the actual Korean corpus converted to implicit NLI task Hendrycks and Gimpel (2017); Sun et al. (2019); Feng et al. (2022), thereby constructing an effective framework coined as Korean ABSA using Pseudo-Classifier with corpus Filtering (KPC-cF) for implicit-feature alignment. Through this process, we assess the impact of our constructed framework on the practical alignment of actual reviews. We validate the pseudo-classifier, generated through the sentence-pair approach, and it outperforms the single approach in translation task. Furthermore, using the model that predicts the translated dataset most effectively as a baseline, we generate pseudo-labels for actual data and conduct real-world testing of Korean ABSA. This involves subsequent fine-tuning the filtered corpus based on language-agnostic embedding similarity for review and aspect sentence pairs, along with setting a threshold for Maximum Softmax Probability (MSP) in pseudo-labels.
@@ 2.1 Motivation and Contribution
The goal of this research is to propose a framework for achieving the best ABSA on actual data with Korean nuances through high-resource languages. Past research by Balahur and Turchi Balahur and Turchi (2012) has shown that machine translation systems can obtain training data for languages other than English in general sentiment classification. However, using translation data for alignment and alignment-free methods Li et al. (2020); Zhang et al. (2021) inadequately address the challenge of universal knowledge transfer for linguistic subgroups like Korean. Also, although it was a different domain at Zhou et al. Zhou et al. (2021), we found it necessary to investigate whether the concept of pseudo-labels could help bridge the gap of feature between translated source data and actual target language data . Therefore, we attempt the following two phases to assess impact of the generated pseudo-classifier, fine-tuned using translated datasets from the ABSA benchmark and pseudo-labeled actual review data, on Korean ABSA. Fig. 2 shows the two-phase pseudo-classifiers we will employ. In the first phase, we investigate the data structure exhibiting performance aligned with by thoroughly learning the fine patterns of the labeled benchmark for each model. NLI structure is referenced based on our research domain and language settings, following the prior work Hoshino et al. (2024). In Phase 2, we fine-tune the baseline , which was effective in training on , by additionally incorporating pseudo-labeled . Employing the tuned model , we conduct predictions and evaluations on manually labeled actual Korean reviews. Throughout this process, Dual filtering are performed to enhance implicit features . Detailed description of our task is provided in Appendix§B.
@@ 2.2 LaBSE based Filtering
In this approach, we aim to extract good-quality sentences-pair from the pseudo-NLI corpus. Language Agnostic BERT Sentence Embedding model Feng et al. (2022) is a multilingual embedding model that supports 109 languages, including some Korean languages. Feng et al. (2022) suggested that the dual-encoder architecture of the LaBSE model, originally designed for machine translation in source-target language data Batheja and Bhattacharyya (2022, 2023), can be applied not only to other monolingual tasks like Semantic Textual Similarity (STS) but also to data (i.e., sentence pair-set) filtering for creating high-quality training corpora in terms of meaning equivalence. Therefore, to mitigate performance degradation caused by the linguistic gap between translated and actual Korean during fine-tuning, we introduce the following filtering method that enables the identification of meaning equivalence or connotation Ghadery et al. (2019) in actual Korean sentence-pairs, even when viewed from the perspective of model trained on bilingual translation pairs. We generate the sentence embeddings for the review text and aspect of the pseudo-NLI corpora using the LaBSE model. Then, we compute the cosine similarity between the review text and aspect sentence embeddings. After that, we extract good quality NLI sentences based on a threshold value of the similarity scores.
@@
We aimed to apply the LaBSE scoring to the actual Korean data , KR3, intending to facilitate flexible learning compared to the translated data , Kor-SemEval (see Appendix§C.1).
| Dataset | Count | LaBSE. S | Pred. Prob |
| (mean) | (mean) | ||
| PL (BaselinemBERT) | 15.23K | 0.22 | 0.986 |
| PL > L0.15 PL > Th0.5 (BaselinemBERT) | 6.77K | 0.24 | 0.988 |
| PL (BaselineXLM-R) | 15.23K | 0.21 | 0.69 |
| PL > L0.15 PL > Th0.5 (BaselineXLM-R) | 6.08K | 0.24 () | 0.74 () |
@@ 2.3 Confidence score Filtering
Meanwhile, we need to develop a classifier capable of optimal predictions on the test, which can be considered as out-of-distribution data separate from the . Drawing on previous research Arora et al. (2021), we expect that language shifts (i.e., ) embody both Background and Semantic shift characteristics. To ensure robust learning in both aspect detection and sentiment classification, we introduce additional thresholding on Maximum Softmax Probability (MSP; Hendrycks and Gimpel 2017) after LaBSE-based filtering on the train set. When considering an input and its corresponding pseudo-label , the score for MSP is expressed as:
Through this, we intended a dual scoring and filtering process to ensure that our does not retrain on misplaced confidence or subpar prediction outcomes for out-of-distribution data. The algorithm for calculate scores and filter with the target batch set can be found in Algorithm 1.
@@ 3.1 Main Experiment: KR3 Test Set
Based on the results from Kor-SemEval, we observed analogous patterns between mBERT and XLM-RBase, notwithstanding their distinct properties (see Appendix§D). Accordingly, We opted for the Baseline-NLI approach (i.e., BaselinemBERT, XLM-R), which demonstrated the best performance, as the base model for Phase 2.
Main Results
To investigate the effect of features for each corpus, we conduct baseline tuning comparisons between the PL and the PL-cF (see Tab. 1, 2 for details). The variants of our tuning framework includes:
Baseline+PL (Pseudo-Labeled data): Fine-tuning Baseline-NLI with pseudo-KR3.
Baseline+PL-cF (Corpus Filtering): Fine-tuning Baseline-NLI with the data obtained by truncating instance from pseudo-KR3, where the threshold of MSP Hendrycks and Gimpel (2017) is less than 0.5 and the cosine similarity between LaBSE embeddings is less than 0.15.
Baseline+TR (TRanslated data)+PL: Fine-tuning Baseline-NLI (pre-tuned or jointly fine-tuned with Kor-SemEval) using pseudo-KR3.
KPC-cF (Baseline+TR+PL-cF): Fine-tuning Baseline-NLI (pre-tuned or jointly fine-tuned with Kor-SemEval) using PL-cF.
Evaluation results on the KR3 test set are presented in Tab. 2 and Fig. 3. We find that the KPC-cF approach achieved adequately trained results in both subtasks for the actual korean data. The model pre-tuned with Kor-SemEval achieves the best performance in Aspect Category Detection (ACD).
For Aspect Category Polarity (ACP), it performs exceptionally well in the tuning of Pseudo-Labels, especially in the Binary setting. Filtered Pseudo-Labels preserve this characteristic well and amplify the performance of all metrics within ACP.
@@ 4 Discussion
In Phase 1, XLM-R, known for its proficiency in capturing cross-lingual representations, exhibits an underfitting tendency concerning the contextual disparities in aspect vocabulary within a single task. This can be attributed to data scarcity relative to model availability for each classifier or viewed as a limitation in single task using SPM in low-resource Korean ABSA Son et al. (2023). Nevertheless, in the NLI task, it showcases potential by outperforming mBERT, guided by "aspect". Conversely, mBERT demonstrates stable results in both single and NLI tasks, exhibiting an overall accuracy increase, particularly in the NLI task (Appendix§D).
Furthermore, Phase 2 reveals that the combination of the NLI approach and translation impacts the metrics of detection in aspects. Pseudo-labels in this phase contribute to enhancing the determination of sentiment, resulting in improved classifier performance. Notably, PL-cF, unlike a mere addition to translated data, play a crucial role in maintaining and enhancing accuracy and F1 score, even with fewer samples. Essentially, the filtered pseudo-NLI set alleviates the bias of pre-trained model parameters during training and enhances performance. This impact of the quality of PL further trained on TR is manifested through the embeddings generated by our model and the ablation results via thresholding (see Fig. 1, 3).
@@
| BaselineKLUE-R | pt | 23.78 | -10.91 |
| BaselineKLUE-R-Large | pt | 13.33 | -18.17 |
| SetFitKLUE-R-13domain | pt, ft | 29.49 | 8.37 |
@@
Looking at Figure 4, In BaselineXLM-R+TR, the Target None class exhibits relatively high softmax probabilities, while data with polarity show relatively low probabilities, indicating significant confidence in polarity and asymmetric learning between classes. The results of KPC-cF demonstrate that additional fine-tuning with filtered data significantly increases the maximum probability of the model for Target Polarity. Simultaneously, with a reduction of over 40% in data, it prevents biased class learning, forming similar distributions among Target classes. Additionally, for polarity data where the target is detected, the regression line shows that higher LaBSE scores correspond to increased confidence. Meanwhile, our KPC-cF approach demonstrated more scalable performance than other training methods in comparisons of intermediate and shallow layer features. KPC-cF demonstrates an improved ability to detect general semantic relationships in Korean, as shown in Tab. 3. In this experiment, we generated zero-shot intermediate embeddings for KorSTS, KLUE-STS Ham et al. (2020); Park et al. (2021) sentence pairs, evaluating similarity with Spearman’s correlation () to Ground Truth. On the other hand, by providing word level features as input to a low-complexity LR model, we further confirmed the efficacy of basic characteristics of filtered data through variations in training outcomes (see Tab. 4). Our framework enhances stability and reliability by optimally thresholding to adjust biased or uncertain data.

### v5

@@ Abstract
Investigations into Aspect-Based Sentiment Analysis (ABSA) for Korean industrial reviews are notably lacking in the existing literature. Our research proposes an intuitive and effective framework for ABSA in low-resource languages such as Korean. It optimizes prediction labels by integrating translated benchmark and unlabeled Korean data. Using a model fine-tuned on translated data, we pseudo-labeled the actual Korean NLI set. Subsequently, we applied LaBSE and MSP-based filtering to this pseudo-NLI set as implicit feature, enhancing Aspect Category Detection and Polarity determination through additional training. Incorporating dual filtering, this model bridged dataset gaps and facilitates feature alignment with minimal resources. By implementing alignment pipelines, our approach aims to leverage high-resource datasets to develop reliable predictive and refined models within corporate or individual communities in low-resource language countries. Compared to English ABSA, our framework showed an approximately 3% difference in F1 scores and accuracy. We will release our dataset and code for Korean ABSA, at this link111https://anonymous.4open.science/r/KPC-cF-21E8.
@@ 1 Introduction
In low-resource downstream tasks such as Korean ABSA, constraints exist in constructing ABSA systems that are socially and industrially beneficial (e.g., obtaining accurate labels and high-quality training data, building a efficient serving model). Addressing this challenge is fundamentally crucial for the practical implementation of multilingual ABSA leveraging the advantages of language models (Zhang et al., 2021; Lin et al., 2023). On the other hand, ABSA utilizing Large Language Models like ChatGPT can perform labeling through prompt tuning. However, due to reduced sensitivity to gradient-based tuning, the model may exhibit overconfidence in random labels, and limitations persist regarding resources for training and inference (Wang et al., 2023; Wu et al., 2023; Kossen et al., 2024).
@@
In this study, we derive pseudo-labels for real Korean reviews using machine-translated English ABSA data, inspired by the past research (Balahur & Turchi, 2012; Hoshino et al., 2024b). Moreover, we employ Dual filtering on the actual Korean corpus converted to implicit NLI task (Hendrycks & Gimpel, 2017; Sun et al., 2019; Feng et al., 2022), thereby constructing an effective framework coined as Korean ABSA using Pseudo-Classifier with Corpus Filtering (KPC-cF) to achieve implicit-feature alignment. Through this process, we assess the impact of our constructed classifier on the practical alignment of actual reviews. We validate that the pseudo-classifier, generated through the sentence-pair approach, outperforms the single approach in translation task. Furthermore, using the model that predicts the translated dataset most effectively as a baseline, we generate pseudo-labels for actual data and conduct real-world testing of Korean ABSA. This involves subsequent fine-tuning the filtered corpus based on language-agnostic embedding similarity for review and aspect sentence pairs, along with setting a threshold for Maximum Softmax Probability (MSP) in pseudo-labels.
@@
2 Task Description
In ABSA, Sun et al. (2019) set the task as equivalent to learning subtasks 3 (Aspect Category Detection) and 4 (Aspect Category Polarity) of SemEval-2014 Task 4 simultaneously. A similar approach was adopted for Korean ABSA in industrial reviews, aiming to develop a task-specific model through comparison of two PLMs (see Appendix§B), differing only in tokenization, vocabulary size, and model parameters (Conneau et al., 2020). Defining a unified-serving model using multi-label-multi-class classification from a task-oriented perspective was considered impractical due to challenges in modifying pre-training set and the ongoing injection of additional data, rendering implicit mapping unattainable (Sun et al., 2020; Ahmed et al., 2022; Qin & Joty, 2022; Kossen et al., 2024). Consequently, problem has been redefined into BERT-based tasks (see Appendix§B.1).
@@ 3.1 Motivation and Contribution
The goal of this research is to propose a framework for achieving the best ABSA on actual data with Korean nuances through high-resource languages. Past research by Balahur & Turchi (2012) has shown that Machine Translation (MT) systems can obtain training data for languages other than English in general sentiment classification. However, Li et al. (2020); Zhang et al. (2021) use explicitly aligned translation features, limiting the exploration of discriminative real target features in low-resource languages like Korean (Fei et al., 2022; Zhao et al., 2024). Also, although it was a different domain at Zhou et al. (2021), we found it necessary to investigate whether the concept of pseudo-labels could help bridge the gap of feature between translated source data and actual target language data . Therefore, we attempted the following two phases to assess the impact of the generated pseudo-classifier, fine-tuned using translated datasets from the ABSA benchmark and pseudo-labeled actual review data, on Korean ABSA. Figure 2 shows the two-phase pseudo-classifiers we will employ. In the first phase, we investigate the data structure exhibiting performance aligned with by thoroughly learning the fine patterns of the labeled benchmark for each model. NLI structure is referenced based on our research domain and language settings, following the prior work (Hoshino et al., 2024a). In Phase 2, we fine-tune the baseline , which was effective in training on , by additionally incorporating pseudo-labeled . Employing the tuned model , we conduct predictions and evaluations on manually labeled actual Korean reviews. Throughout this process, Dual filtering is performed to enhance implicit features . Detailed description of our training strategy is provided in Section§3.4, Appendix§C.
@@ 3.2 LaBSE based Filtering
Our main approach, we aim to extract good-quality sentences-pair from the pseudo-NLI corpus. Language Agnostic BERT Sentence Embedding model (Feng et al., 2022) is a multilingual embedding model that supports 109 languages, including some Korean languages. Feng et al. (2022) suggested that the dual-encoder architecture of the LaBSE model, originally designed for machine translation in source-target language data (Batheja & Bhattacharyya, 2022, 2023), can be applied not only to other monolingual tasks like Semantic Textual Similarity (STS) but also to data (i.e., sentence pair-set) filtering for creating high-quality training corpora in terms of meaning equivalence. Therefore, to mitigate performance degradation caused by the linguistic gap between translated and actual Korean during fine-tuning, we introduce the following filtering method that enables the identification of meaning equivalence or connotation (Ghadery et al., 2019) in actual Korean sentence-pairs, even when viewed from the perspective of model trained on bilingual translation pairs. We generate the sentence embeddings for the review text and aspect of the pseudo-NLI corpora using the LaBSE model. Then, we compute the cosine similarity between the review text and aspect sentence embeddings. After that, we extract good quality NLI sentences based on a threshold value of the similarity scores.
@@
We aimed to apply the LaBSE scoring to the actual Korean dataset , KR3, intending to facilitate flexible learning compared to the Eng-Kor translated dataset , Kor-SemEval (see Section§4.1).
@@ 3.3 Confidence score Filtering
Meanwhile, we need to develop a classifier capable of optimal predictions on the test, which can be considered as out-of-distribution data separate from the . Drawing on previous research (Arora et al., 2021), we expect that language shifts (i.e., ) embody both Background and Semantic shift characteristics. To ensure robust learning in both aspect detection and sentiment classification, we introduce additional thresholding on Maximum Softmax Probability (MSP; Hendrycks & Gimpel 2017) after LaBSE-based filtering on the train set. When considering an input and its corresponding pseudo-label , the score for MSP is expressed as:
Through this, we intended a dual scoring and filtering process to ensure that our does not retrain on misplaced confidence or subpar prediction outcomes for out-of-distribution data. This process is defined by the following expected success rate of the trained models, where filters the sentence pairs and pseudo-labels.
Our algorithm for calculate scores and filter with the target batch set can be found in Algorithm 1.
3.4 Target-aligned Objective
In this section, we focus on the langauge adaptation of ABSA. The input data includes a set of labeled sentences from a translated source data and a set of unlabeled sentences from a target language data . Our goal is to validate our filtering technique across three types of tasks: (1) training only pseudo-labeled , (2) joint training using both and , (3) transfer learning from to . Namely, unlike Chen et al. (2023), we maintain token embeddings and fine-tuned the transformer body & head. Contrary to Gururangan et al. (2020), we trained with and then reinitialized the classifier layer for further fine-tuning. This approach examines if the universal encoder and params aligned with enhance implicit alignment via pseudo-labels of refined .
was also evaluated, despite its inefficiency in terms of memory and computation, for data validation purposes. Thus, shuffling was used in .
Assumption 3.1 (Contextual, Sementical gradient).
Assumption 3.2 (Target-Aligned Feature).
Under Assumptions in objective, for any example , after the gradient step at timestep , the value embedding of the sentence-pair tokens is more predictive of the label .
4 Experimental Setup
4.1 Dataset for Fine-Tuning and Test
Kor-SemEval
We translate the SemEval-2014 Task 4 (Pontiki et al., 2014) dataset333http://alt.qcri.org/semeval2014/task4/. Moreover, it is evaluated for Korean ABSA. The training data was machine-translated (by Google Translate), and Test data was corrected manually only for fewer than 10 instances where abnormal translations occurred after machine translation. Each sentence contains a list of aspect with the sentiment polarity . Ultimately, given a sentence in the sentence, we need to:
detects the mention of an aspect ;
This setting allows us to jointly evaluate Subtask 3 (Aspect Category Detection) and 4 (Aspect Category Polarity).
@@ 4.1 Dataset for Fine-Tuning and Test
| GPT-4o mini (OpenAI, 2024) | - - | - | 28.00 | 99.91* | 43.74 | 84.41 | 86.13 | 89.10 | |
| Zero-Shot (Li et al., 2020) | 1.43MB 0.38M | 75.31 | 74.41 | 74.15 | 81.52 | 84.00 | 88.73 | ||
| Trans-Ta (Zhang et al., 2021) | 1.54MB 0.37M | 94.04* | 78.94 | 85.78 | 74.84 | 77.37 | 80.75 | ||
@@
KR3
Unlike the domains previously used for Korean sentiment classification (Lee et al., 2020; Yang, 2021; Ban, 2022), Korean Restaurant Review with Ratings (KR3) is a comprehensive dataset encompassing various food service establishments, constructed from actual certified map reviews. In the case of restaurant reviews, words and expressions that evaluate positive and negative are mainly included, and real users often infer what a restaurant is like by looking at its reviews. Accordingly, Jung et al.3 constructed the KR3 dataset by crawling and preprocessing user reviews and star ratings of websites that collect broad food service information and ratings. KR3 has 388,111 positive and 70,910 negative, providing a total of 459,021 data plus 182,741 unclassified data, and distributed to Hugging Face444https://huggingface.co/datasets/leey4n/KR3. We structured our training and test datasets to match the size of Kor-SemEval. Specifically, we addressed potential biases by randomly sampling indices from the original KR3, ensuring that evaluations for a specific restaurant were non-overlapping. Additionally, we maintained an even distribution of positive, negative, and neutral (ambiguous) classes, irrespective of the aspects indicated in the original KR3. This preprocessing step aimed to capture a comprehensive representation of sentiments across diverse attributes of sentences in the dataset. Subsequently, the data were configured to suit sentence pair classification. To allocate polarity labels for each aspect within the KR3 dataset, pseudo-labeling was conducted utilizing the optimal model identified during the Kor-SemEval performance evaluation. Pseudo-labels were assigned to the KR3 training data, and post pseudo-labeling, the test data underwent manual re-labeling by researchers.
Table 8 shows some Kor-SemEval and KR3 training data samples. In the case of KR3, the negative aspect is better reflected. Meanwhile, while Kor-SemEval gave neutrality to mediocre service, KR3 did not give neutrality to mediocre taste. While positive and negative data have been sufficiently accumulated and reflected, the tendency for a lack of neutral data can be confirmed in advance through some samples. We have organized both Kor-SemEval and KR3 data as open-source to facilitate their use in various training and evaluation scenarios.
@@ 5.1 Main Experiment: KR3 Test Set
On the other hand, based on the results from Kor-SemEval, we examined the dissimilarity specific to the trans-align task between mBERT and XLM-RBase (see Appendix§D). Accordingly, We opted for the Baseline-NLI approach (i.e., BaselinemBERT, XLM-R), which demonstrated the best performance, as the base model for Phase 2 (see Figure 2, Section§3.1). We conducted evaluations on KR3 test data.
5.2 Main Results
To investigate the effect of features for each corpus, we conduct baseline tuning comparisons between the PL and the PL-cF (see Table 1, 2 for details). The variants of our tuning framework includes:
(1) BaselinePLM+PL (Pseudo-Labeled data): Fine-tuning the Baseline PLM in an NLI manner with pseudo-KR3.
(2) BaselinePLM+PL-cF (Corpus Filtering): Fine-tuning Baseline PLM with the NLI data obtained by rejecting instance from pseudo-KR3, where the threshold of MSP (Hendrycks & Gimpel, 2017) is less than 0.5 and the cosine similarity between LaBSE embeddings is less than 0.15.
(3) BaselinePLM+TR (TRanslated data)+PL: Fine-tuning Baseline PLM in an NLI manner (pre-tuned or jointly fine-tuned with Kor-SemEval) using pseudo-KR3.
(4) KPC-cF (BaselinePLM+TR+PL-cF): Fine-tuning Baseline PLM in an NLI manner (pre-tuned or jointly fine-tuned with Kor-SemEval) using PL-cF.
Results on the KR3 test set are presented in Table 1 and Figure 5. We find that the KPC-cF approach achieved adequately trained results in both subtasks for the actual korean data. The model pre-tuned with Kor-SemEval achieves the best performance in Aspect Category Detection (ACD). For Aspect Category Polarity (ACP), it performs exceptionally well in the tuning of Pseudo-Labels, especially in the Binary setting. Filtered Pseudo-Labels preserve this characteristic well and amplify the performance of all metrics within ACP.
| Dataset | Count | LaBSE. S | Pred. Prob | Corr.C |
| (mean) | (mean) | (target none) | ||
| PL (BaselinemBERT) | 15.23K | 0.22 | 0.986 | -0.08 |
| PL PL (BaselinemBERT) | 6.77K | 0.24 | 0.988 | -0.05 |
| PL (BaselineXLM-R) | 15.23K | 0.21 | 0.69 | -0.09 |
| PL PL (BaselineXLM-R) | 6.08K | 0.24 | 0.74 | -0.12 |
@@ 6 Discussion
In Phase 1, XLM-R, known for its proficiency in capturing cross-lingual representations, exhibits an underfitting tendency concerning the contextual disparities in aspect vocabulary within a single task. This can be attributed to data scarcity relative to model availability for each classifier or viewed as a limitation in single task using SPM in low-resource Korean ABSA (Son et al., 2023). Nevertheless, in the NLI task, it showcases potential by outperforming mBERT, guided by ”aspect”. Conversely, mBERT demonstrates stable results in both single and NLI tasks, exhibiting an overall accuracy increase, particularly in the NLI task (see Appendix§D). Furthermore, the pseudo-labels (PL) generated in Phase 2 refine polarity estimation, thereby enhancing classifier performance. Notably, PL-cF plays a crucial role in maintaining and improving accuracy and F1 score, even with fewer samples, unlike a simple addition of PL. The impact of further training on PL (PL further trained on TR) is evident through embedding analysis and ablation studies based on thresholding (see Figure 1, 5). Additionally, GPT-4o mini exhibits high recall but significantly lower precision, as it assigns polarity-labeled aspects to a much larger set. In contrast, Trans-Ta achieves higher precision, suggesting that alignment models trained on translated data can refine contextual categories relevant to domain classification. This further implies the importance of conservative evaluation in review classification, particularly for aspects that may impact reputation, as well as its significance in real-time inference.
@@
| BaselineKLUE-R | pre-trained | 23.78 | -10.91 |
| BaselineKLUE-R-Large | pre-trained | 13.33 | -18.17 |
| SetFitKLUE-R-13domain | pre & fine-tuned | 29.49 | 8.37 |
@@
Looking at Figure 3, compared to BaselineXLM-R+TR, KPC-cF significantly increases the maximum probability of Target Polarity through additional fine-tuning with filtered data. Despite a 40% data reduction, it prevents biased class learning and maintains a balanced probability distribution across Target classes. Furthermore, for sentiment polarity data where the target is detected, higher LaBSE scores correspond to increased model confidence. KPC-cF also exhibits scalability in intermediate and shallow-layer feature comparisons. It enhances the model’s ability to capture general contextual relationships between sentence embeddings (see Table 3), as confirmed through zero-shot embedding evaluations using correlation coefficients against Ground Truth. Additionally, training a linear model with word-level features further validates the influence of filtered data characteristics on training outcomes (see Table 4). This confirms effectiveness of our approach in refining semantic label information. Our framework improves model stability and reliability by optimally rejecting uncertain or biased data.